# Zero-shot Musical Stem Retrieval with Joint-Embedding Predictive Architectures

### v1

@@ II Method
Figure 1: Overview of our model. In both phases, it is trained using a pair composed of a context mix $\mathbb{x}$ and a target stem $\mathbb{\bar{x}}$ extracted from the same track. In Phase 1 (left), we pretrain only the encoder $f_{\theta}$ using contrastive learning to bring together the (averaged) representations $\mathbb{s}$ and $\mathbb{\bar{s}}$ of the context and target. In Phase 2 (right), a predictor $g_{\phi}$, conditioned on the instrument of the target stem $c$, tries to retrieve the patchwise representations of the target stem $\mathbb{\bar{z}}$ from the ones of the context mix $\mathbb{z}$. In this phase, the parameters of the target encoder $f_{\bar{\theta}}$ are updated as an EMA of the ones of the encoder $f_{\theta}$.
@@ II-A Phase 1: Contrastive pretraining
First, we pretrain only the encoder $f_{\theta}$ using contrastive learning. Given a chunk of audio composed of $S$ stems, we first pick one as the target $\mathbb{\bar{x}}$ and a random subset of the remaining ones as the context $\mathbb{x}$. $\mathbb{x}$ and $\mathbb{\bar{x}}$ are then turned into LMS, divided into a regular time and frequency grid of $K$ non-overlapping patches, and fed into the encoder $f_{\theta}$. This encoder returns sequences of patch-wise embeddings $\mathbb{z},\mathbb{\bar{z}}\in\mathbb{R}^{K\times d}$, with $d$ being the latent dimension.
We then use these embeddings’ mean over $K$ as the latent representations $\mathbb{s},\mathbb{\bar{s}}\in\mathbb{R}^{d}$ of the context mix and target stem, respectively.
Finally, the parameters $\theta$ of the encoder are optimized by minimizing the contrastive loss $\mathcal{L}_{c}$ within batches of representations, using $(\mathbb{s},\mathbb{\bar{s}})$ as a positive pair and the other elements of the batch $B$ as negative samples, as in [4]:
$\mathcal{L}_{c}(\mathbb{s},\mathbb{\bar{s}})=-\log\left(\dfrac{\exp(\text{sim}(\mathbb{s},\mathbb{\bar{s}})/\tau)}{\sum_{\mathbb{s}^{\prime}\in B\setminus\{\mathbb{s}\}}\exp(\text{sim}(\mathbb{s},\mathbb{s}^{\prime})/\tau)}\right),$ (1)
@@ II-B Phase 2: Joint-Embedding Predictive Architecture
The second phase is similar to [2]. We randomly pick a context mix $\mathbb{x}$, a target stem $\mathbb{\bar{x}}$, and turn them into LMS. We denote the instrument label of the target stem as $c$. The context mix passes through the encoder $f_{\theta}$, which returns a sequence of embeddings $\mathbb{z}=(\mathbb{z}_{1},\dots,\mathbb{z}_{K})\in\mathbb{R}^{K\times d}$. Conversely, we obtain $\mathbb{\bar{z}}=(\mathbb{\bar{z}}_{1},\dots,\mathbb{\bar{z}}_{K})\in\mathbb{R}^{K\times d}$ by passing the target stem $\mathbb{\bar{x}}$ through a target encoder $f_{\bar{\theta}}$ with parameters $\bar{\theta}$.
Then, a predictor $g_{\phi}$, conditioned on the target instrument label $c$, produces predictions $\mathbb{\tilde{z}}_{i}\in\mathbb{R}^{d}$ from the context embeddings, i.e., for all $k\in\{1,\dots,K\}$:
$\mathbb{\tilde{z}}_{k}=\ g_{\phi}(\mathbb{z}_{k},c).$ (2)
@@
$\mathcal{L}(\mathbb{\tilde{z}},\mathbb{\bar{z}})=\sum_{k=1}^{K}\left\|\dfrac{\mathbb{\tilde{z}}_{k}}{\|\mathbb{\tilde{z}}_{k}\|}-\dfrac{\mathbb{\bar{z}}_{k}}{\|\mathbb{\bar{z}}_{k}\|}\right\|^{2}.$ (3)
@@ II-C Predictor conditioning
Let $\mathbb{c}=\text{CLAP}(c)\in\mathbb{R}^{p}$ be the embedding used for conditioning the predictor. While in [2], this instrument class embedding is concatenated to the context embeddings $\mathbb{z}_{i}$, here we instead use FiLM conditioning [14]. More precisely, our predictor $g_{\phi}$ is a MLP with $L$ layers and $m$ hidden units. For $l\in\{1,\dots,L-1\}$, we learn two affine mappings $\beta_{l}$ and $\gamma_{l}:\mathbb{R}^{p}\to\mathbb{R}^{m}$. Let $\mathbb{h}_{l}$ be the output of the $l$-th layer of $g_{\phi}$, the input of the $l+1$-th layer is:
$\widehat{\mathbb{h}}_{l+1}=\text{ReLU}\left(\gamma_{l}(\mathbb{c})\cdot\mathbb{h}_{l}+\beta_{l}(\mathbb{c})\right),$ (5)
@@ III-A Experimental setup
We test the ability of our model to retrieve the missing stem from a given mix when conditioned on the instrument label of the missing stem. Our dataset comprises $N$ tracks $\mathbb{x}^{(1)},\dots,\mathbb{x}^{(N)}$, and each track $\mathbb{x}^{(n)}$ is composed of $S_{n}$ stems $\mathbb{x}^{(n)}_{1},\dots,\mathbb{x}^{(n)}_{S_{n}}$.
Given an individual stem $\mathbb{x}^{(n)}_{s}$, $\mathbb{x}^{(n)}_{\neg s}$ denotes the mix containing all stems from $\mathbb{x}^{(n)}$ without $\mathbb{x}^{(n)}_{s}$. We first encode all individual stems and average the resulting embeddings over the patch dimensions to obtain a reference set =$\mathbb{Z}=\{\mathbb{z}^{(n)}_{s}\}$ of latent representations.
In addition, we pass the mixes $\mathbb{x}^{(n)}_{\neg s}$ through the encoder, whose outputs are fed into the predictor conditioned on the corresponding instrument label of $\mathbb{x}^{(n)}_{s}$ to produce queries $\mathbb{q}^{(n)}_{s}$. The queries are averaged similarly to the representations $\mathbb{z}^{(n)}_{s}$. For such a query $\mathbb{q}^{(n)}_{s}$, we then measure the distance to the representation of the corresponding stem $\mathbb{z}^{(n)}_{s}$. To do so, we rely on two metrics:
  • The Recall at $k$ measures the proportions of queries $\mathbb{q}^{(n)}_{s}$ for which the missing stem $\mathbb{z}^{(n)}_{s}$ lies in their $k$ nearest neighbors. We report the recall at $k$ of the different models for $k\in\{1,5,10\}$.
  • The Normalized Rank [10] of a query $\mathbb{q}^{(n)}_{s}$ is the position of the ground-truth stem $\mathbb{z}^{(n)}_{s}$ in the sorted list of distances $\{\|\mathbb{q}^{(n)}_{s}-\mathbb{z}\|\}_{\mathbb{z}\in\mathbb{Z}}$, normalized by the total number of stems $|\mathbb{Z}|$. For each model, we report the mean and median Normalized Ranks over all queries.
@@ III-D Instrument-specific analysis
IV Temporal information in embeddings
@@
We concatenate the outputs of the encoder along the frequency axis. Each embedding encodes a 160 ms segment of the audio signal and is up-sampled using $K=8$ linear probes. Each probe targets a 20 ms sub-segment within the 160 ms window and outputs a beat activation function. We use 8-fold cross-validation [^3] We train on Ballroom [20], Hainsworth [21], SMC [22], and Harmonix [23] and keep GTZAN [24] reserved as an unseen test set. and report F-measure, AMLt, and CMLt, following standard protocols [25, 26]. The probes are trained with binary cross-entropy loss and label widening [26].
To predict beat positions, the activation is fed into a dynamic Bayesian network [27]. Results for GTZAN and SMC are presented in Table II. We compare our performances with a standard supervised method [26], as well as another approach that also consists in SSL pre-training followed by linear probing [28].
TABLE II: Beat-tracking performance in 8-fold cross-validation setting.
@@
Our model achieves decent results, with AMLt scores close to those of the baselines. Even if not as good as the other methods, especially in terms of F-measure and CMLt, these results reveal that, despite a patch resolution of only 160 ms, the embeddings learned by our model retain precise temporal information. This makes our approach promising for more fine-grained tasks.
V Conclusion
In this paper, we introduce the first model for zero-shot musical stem retrieval using JEPAs. We demonstrate the effectiveness of our design choices by evaluating our model on different benchmarks, and we significantly outperform previous baselines. In addition, as our results on beat tracking reveal, our embeddings are not only useful for musical stem retrieval but also retain temporal information, making them potentially usable for both global and local MIR tasks or music accompaniment generation.

### v2

@@ Abstract
Disclaimer: This paper is an extended version of our ICASSP 2025 publication. The ICASSP version is available at https://arxiv.org/pdf/2411.19806v1.
@@ II Method
Figure 1: Overview of our model. In both phases, it is trained using a pair composed of a context mix $\mathbf{x}$ and a target stem $\mathbf{\bar{x}}$ extracted from the same track. In Phase 1 (left), we pretrain only the encoder $f_{\theta}$ using contrastive learning to bring together the (averaged) representations $\mathbf{s}$ and $\mathbf{\bar{s}}$ of the context and target. In Phase 2 (right), a predictor $g_{\phi}$, conditioned on the instrument of the target stem $c$, tries to retrieve the patchwise representations of the target stem $\mathbf{\bar{z}}$ from the ones of the context mix $\mathbf{z}$. In this phase, the parameters of the target encoder $f_{\bar{\theta}}$ are updated as an EMA of the ones of the encoder $f_{\theta}$.
@@ II-A Phase 1: Contrastive pretraining
First, we pretrain only the encoder $f_{\theta}$ using contrastive learning. Given a chunk of audio composed of $S$ stems, we first pick one as the target $\mathbf{\bar{x}}$ and a random subset of the remaining ones as the context $\mathbf{x}$. $\mathbf{x}$ and $\mathbf{\bar{x}}$ are then turned into LMS, divided into a regular time and frequency grid of $K$ non-overlapping patches, and fed into the encoder $f_{\theta}$. This encoder returns sequences of patch-wise embeddings $\mathbf{z},\mathbf{\bar{z}}\in\mathbb{R}^{K\times d}$, with $d$ being the latent dimension.
We then use these embeddings’ mean over $K$ as the latent representations $\mathbf{s},\mathbf{\bar{s}}\in\mathbb{R}^{d}$ of the context mix and target stem, respectively.
Finally, the parameters $\theta$ of the encoder are optimized by minimizing the contrastive loss $\mathcal{L}_{c}$ within batches of representations, using $(\mathbf{s},\mathbf{\bar{s}})$ as a positive pair and the other elements of the batch $B$ as negative samples, as in [4]:
$\mathcal{L}_{c}(\mathbf{s},\mathbf{\bar{s}})=-\log\left(\dfrac{\exp(\text{sim}(\mathbf{s},\mathbf{\bar{s}})/\tau)}{\sum_{\mathbf{s}^{\prime}\in B\setminus\{\mathbf{s}\}}\exp(\text{sim}(\mathbf{s},\mathbf{s}^{\prime})/\tau)}\right),$ (1)
@@ II-B Phase 2: Joint-Embedding Predictive Architecture
The second phase is similar to [2]. We randomly pick a context mix $\mathbf{x}$, a target stem $\mathbf{\bar{x}}$, and turn them into LMS. We denote the instrument label of the target stem as $c$. The context mix passes through the encoder $f_{\theta}$, which returns a sequence of embeddings $\mathbf{z}=(\mathbf{z}_{1},\dots,\mathbf{z}_{K})\in\mathbb{R}^{K\times d}$. Conversely, we obtain $\mathbf{\bar{z}}=(\mathbf{\bar{z}}_{1},\dots,\mathbf{\bar{z}}_{K})\in\mathbb{R}^{K\times d}$ by passing the target stem $\mathbf{\bar{x}}$ through a target encoder $f_{\bar{\theta}}$ with parameters $\bar{\theta}$.
Then, a predictor $g_{\phi}$, conditioned on the target instrument label $c$, produces predictions $\mathbf{\tilde{z}}_{i}\in\mathbb{R}^{d}$ from the context embeddings, i.e., for all $k\in\{1,\dots,K\}$:
$\mathbf{\tilde{z}}_{k}=\ g_{\phi}(\mathbf{z}_{k},c).$ (2)
@@
$\mathcal{L}(\mathbf{\tilde{z}},\mathbf{\bar{z}})=\sum_{k=1}^{K}\left\|\dfrac{\mathbf{\tilde{z}}_{k}}{\|\mathbf{\tilde{z}}_{k}\|}-\dfrac{\mathbf{\bar{z}}_{k}}{\|\mathbf{\bar{z}}_{k}\|}\right\|^{2}.$ (3)
@@ II-C Predictor conditioning
Let $\mathbf{c}=\text{CLAP}(c)\in\mathbb{R}^{p}$ be the embedding used for conditioning the predictor. While in [2], this instrument class embedding is concatenated to the context embeddings $\mathbf{z}_{i}$, here we instead use FiLM conditioning [14]. More precisely, our predictor $g_{\phi}$ is a MLP with $L$ layers and $m$ hidden units. For $l\in\{1,\dots,L-1\}$, we learn two affine mappings $\beta_{l}$ and $\gamma_{l}:\mathbb{R}^{p}\to\mathbb{R}^{m}$. Let $\mathbf{h}_{l}$ be the output of the $l$-th layer of $g_{\phi}$, the input of the $l+1$-th layer is:
$\widehat{\mathbf{h}}_{l+1}=\text{ReLU}\left(\gamma_{l}(\mathbf{c})\cdot\mathbf{h}_{l}+\beta_{l}(\mathbf{c})\right),$ (5)
@@ III-A Experimental setup
We test the ability of our model to retrieve the missing stem from a given mix when conditioned on the instrument label of the missing stem. Our dataset comprises $N$ tracks $\mathbf{x}^{(1)},\dots,\mathbf{x}^{(N)}$, and each track $\mathbf{x}^{(n)}$ is composed of $S_{n}$ stems $\mathbf{x}^{(n)}_{1},\dots,\mathbf{x}^{(n)}_{S_{n}}$.
Given an individual stem $\mathbf{x}^{(n)}_{s}$, $\mathbf{x}^{(n)}_{\neg s}$ denotes the mix containing all stems from $\mathbf{x}^{(n)}$ but $\mathbf{x}^{(n)}_{s}$. We first encode all individual stems and average the resulting embeddings over the patch dimensions to obtain a reference set $\mathbf{Z}=\{\mathbf{z}^{(n)}_{s}\}$ of latent representations.
In addition, we pass the mixes $\mathbf{x}^{(n)}_{\neg s}$ through the encoder, whose outputs are fed into the predictor conditioned on the corresponding instrument label of $\mathbf{x}^{(n)}_{s}$ to produce queries $\mathbf{q}^{(n)}_{s}$. The queries are averaged similarly to the representations $\mathbf{z}^{(n)}_{s}$. For such a query $\mathbf{q}^{(n)}_{s}$, we then measure the distance to the representation of the corresponding stem $\mathbf{z}^{(n)}_{s}$. To do so, we rely on two metrics:
  • The Recall at $k$ measures the proportions of queries $\mathbf{q}^{(n)}_{s}$ for which the missing stem $\mathbf{z}^{(n)}_{s}$ lies in their $k$ nearest neighbors. We report the recall at $k$ of the different models for $k\in\{1,5,10\}$.
  • The Normalized Rank [10] of a query $\mathbf{q}^{(n)}_{s}$ is the position of the ground-truth stem $\mathbf{z}^{(n)}_{s}$ in the sorted list of distances $\{\|\mathbf{q}^{(n)}_{s}-\mathbf{z}\|\}_{\mathbf{z}\in\mathbf{Z}}$, normalized by the total number of stems $|\mathbf{Z}|$. For each model, we report the mean and median Normalized Ranks over all queries.
@@ III-D Instrument-specific analysis
IV Music representation learning
In this section, we investigate the musical features encoded in the representations learned by our model. We hypothesize that the encoder captures shared musical information among different stems of the same track, such as rhythm or harmony, to aid the predictor. To verify this, we evaluate it on several downstream classification tasks, a standard protocol for representation learning methods [20, 13, 21, 22].
Specifically, each audio sample from the downstream dataset is processed by the encoder, with its patch-wise outputs concatenated and averaged along the frequency and time dimensions to obtain a single 3840-dimensional global embedding per audio, as done in [2, 13].
These embeddings are passed through an MLP with 512 hidden units and a softmax layer, which is trained by minimizing the cross-entropy between the predicted distribution and the ground truth labels. To limit the bias induced by the optimization of the probe, results are computed as the average performance of three probes with identical hyperparameters but different seeds. In addition, we try different learning rates and report the best results, as described in the constrained evaluation track of [22].
IV-A Downstream tasks
To validate our hypothesis, we focus on global musical features that are shared among the different stems of a track, namely tagging, key, and genre estimation. Additionally, we include an instrument classification task to observe whether the encoder preserves stem-specific information. For facilitating comparisons to existing work, we also pick our downstream tasks from the MARBLE benchmark [22]. The full list of datasets and associated tasks is depicted in Table II. For a more in-depth description of the datasets, tasks, and corresponding metrics, we refer the reader to [22].
TABLE II: Datasets used for downstream tasks.
Dataset classes Task
Giantsteps (GS) [23] 24 Key detection
GTZAN [24] 10 Genre classification
MagnaTagATune (MTT) [25] 50 Auto-tagging
NSynth [26] 11 Instrument classification
IV-B Results
Compared to the original Stem-JEPA restricted to 4 stems, we observe overall better performances. Conditioning the predictor with CLAP embeddings consistently improves performances on all downstream tasks.
Contrastive pretraining, while effective for key-related tasks, leads to performance degradation on other tasks. For example, on GTZAN (genre classification) and MTT (auto-tagging), JEPA models with contrastive pretraining underperform compared to their standalone JEPA or contrastive baseline counterparts. This suggests that our naive contrastive pretraining strategy—using identical hyperparameters and training durations for both contrastive and JEPA phases—may not be optimal. Adjustments such as fewer pretraining steps or higher contrastive loss temperatures could potentially yield better results.
Finally, conditioning with FiLM layers improves performances for all tasks.
TABLE III: Influence of our different design choices on various downstream tasks, and comparison with existing baselines. Best results are bold and second-to-best are underlined.
GS GTZAN MTT NSynth
Model Accrefined Acc ROC AP Acc
Contrastive baseline 66.4 68.9 89.4 41.6 76.8
Stem-JEPA [2] 40.2 68.6 89.9 42.8 73.5
+CLAP conditioning 50.0 76.7 90.5 44.0 75.5
+contrastive pretraining 58.7 55.6 88.0 37.7 74.5
+FiLM conditioning 60.0 63.0 88.8 39.8 76.7
MULE [20] 64.9 75.5 91.2 40.1 74.6
Jukebox [27] 63.8 77.9 91.4 40.6 70.4
IV-C Comparison with baselines.
Our models outperform MULE and Jukebox on most tasks, except auto-tagging on the MTT dataset. However, none of our models consistently surpass these baselines across all tasks.
It is worth noting that both MULE and Jukebox were trained on significantly larger datasets—117k hours for MULE and 60k–120k hours for Jukebox—while our models were trained on just 1,350 hours of data, roughly 100 times less. This suggests that leveraging separated sources for music representation learning is a promising direction for reducing data requirements.
V Temporal information in embeddings
@@
V-A Experimental setup
[…]
To predict beat positions, the activation is fed into a dynamic Bayesian network [34].
We compare our method against two baseline approaches: a supervised beat tracking model [33], optimized specifically for this task and a self-supervised approach using contrastive pretraining followed by linear probing [35], evaluated under the same protocol as our method.
V-B Results
TABLE IV: Beat-tracking performance in 8-fold cross-validation setting.
@@ V-B Results
Table IV presents the results on GTZAN and SMC datasets. Our model achieved reasonable results, particularly in terms of AMLt, where scores were close to those of the baselines. However, the model underperformed compared to the baselines in terms of F-measure and CMLt.
[…]
VI Conclusion
In this paper, we introduce the first model for zero-shot musical stem retrieval using JEPAs. We build upon the work of [2] and extend it to arbitrary instruments, improving its usability in real-world scenarios. Additionally, we explore alternative design choices for the predictor, replacing concatenation with FiLM conditioning, and propose pre-training the encoder via contrastive learning.
Our results indicate that these modifications not only generalize Stem-JEPA but also enhance its performance, both on retrieval tasks even in the 4-stem configuration (see Table I), and across various MIR downstream tasks, occasionally outperforming top-performing baselines (see Table III).
However, our experiments reveal the challenge of consistently achieving the best of both worlds. Across multiple tasks, there is no single design choice that emerges as clearly superior. This suggests that further research is necessary to identify an optimal balance, both in terms of pre-training strategies and predictor conditioning methods.
In addition, as our results on beat tracking reveal, our embeddings are not only useful for musical stem retrieval but also retain temporal information, making them potentially usable for both global and local MIR tasks or music accompaniment generation.
@@ VII Acknowledgments
We would like to thank Amaury Delort for his support regarding data management and preprocessing.